Consider the following jobs having I/0 wait of 70% with Arrival time and CPU minutes needed.

Jobs Arrival Time CPU minutes needed 13:10 8 2 13:30 6 3 13:40 4 4 13:50 4 5 13:55 3

Calculate CPU idle, CPU busy and CPU busy/ process. Draw a graph showing no. of jobs and when it completes.
Answer #1

Here, based on the given task set, we task are scheduled on the basis of their arrival i.e first come first server basis.

  1. Since all arrive at 13:XX , the 13 until of time is related, hence the arrival time is taken with respect to 13 i.e arrival time will be 10, 30, 40, 50, and 55 units of time.
  2. Based on this data , the Graph when the jobs are arrived and completed is given below :5 6 8 9 11 12 18 19 10 A 1/8 A 13 A 4/8 14 А 5/8 15 А 6/8 16 A 718 17 А. 8/8 0/0 0/0 0/0 0/0 0/0 0/0 0/0 0/0 0/0 0/0 18 3/8 0

    Based on this graph , the

    CPU idle

    = Number of empty slots in the graph / Total number of slots in the graph =  

    = 34 / 59 = 57.62 %

    CPU busy

    = Number of filled slots in the graph / Total number of slots in the graph

    = 25 / 59 = 42.38 %

    CPU busy/ process

    =  Number of filled slots in the graph / Number of process scheduled

    = 25 / 5 = 5